# Greener GRASS: Enhancing GNNs with Encoding, Rewiring, and Attention

### v4

@@ 3.3 Attention Mechanism
We use post-normalization in residual connections, which has been shown to improve the expressiveness of Transformers Liu et al. (2020). The residual connection is scaled by the constant $\alpha$ to improve training stability Wang et al. (2022).
@@ F.2 Hyperparameters
Model ZINC MNIST CIFAR10 PATTERN CLUSTER # Parameters 496545 103690 103738 495298 495558 # Attention Layers 49 15 15 53 53 Attention Layer Dim. 32 24 24 32 32 Task Head Hidden Dim. 192 144 144 N/A (Linear) N/A (Linear) # Epochs 2000 200 400 500 50 Warmup Epoch Ratio 0.1 0.05 0.1 0.1 0.1 Batch Size 200 200 200 200 200 Initial Learning Rate 1e-7 1e-7 1e-7 1e-7 1e-7 Peak Learning Rate 5e-4 1e-3 1e-3 1e-3 1e-3 Final Learning Rate 1e-7 1e-7 1e-7 3e-4 1e-7 Betas (0.95, 0.98) (0.95, 0.98) (0.95, 0.98) (0.95, 0.98) (0.95, 0.98) Weight Decay Factor 0.5 0.3 0.3 3.0 0.3 Label Smoothing Factor N/A (Regression) 0.1 0.1 0.1 0.1 Random Walk Encoding Type RRWP RRWP RRWP RRWP RRWP (D-)RRWP Random Walk Length 32 24 24 32 32 Random Regular Graph Degree 6 6 6 6 6 Random Edge Removal Rate 0.1 0.1 0.1 0.5 0.5
@@
Model Peptides-func Peptides-struct PascalVOC-SP COCO-SP roman-empire # Parameters 500074 498315 501493 499377 2075730 # Attention Layers 48 48 53 53 24 Attention Layer Dim. 32 32 32 32 96 Task Head Hidden Dim. 192 192 N/A (Linear) N/A (Linear) N/A (Linear) # Epochs 500 500 500 100 4000 Warmup Epoch Ratio 0.1 0.1 0.1 0.1 0.1 Batch Size 200 200 200 200 1* Initial Learning Rate 1e-7 1e-7 1e-7 1e-7 1e-7 Peak Learning Rate 1e-3 1e-3 1e-3 1e-3 1e-3 Final Learning Rate 1e-7 1e-7 1e-7 1e-7 1e-7 Betas (0.95, 0.98) (0.95, 0.98) (0.95, 0.98) (0.95, 0.98) (0.95, 0.98) Weight Decay Factor 0.3 3.0 1.0 0.3 1.0 Label Smoothing Factor 0.1 N/A (Regression) 0.1 0.1 0.1 Random Walk Encoding Type D-RRWP RRWP D-RRWP D-RRWP D-RRWP (D-)RRWP Random Walk Length 128 64 128 64 256 Random Regular Graph Degree 3 3 6 6 3 Random Edge Removal Rate 0.1 0.1 0.1 0.1 0.5

### v5

@@ 3.3 Attention Mechanism
We use post-normalization in residual connections, which has been shown to improve the expressiveness of Transformers Liu et al. (2020). The residual connection is scaled by a constant $\alpha$ to improve training stability Wang et al. (2022).
@@ F.2 Hyperparameters
Model ZINC MNIST CIFAR10 PATTERN CLUSTER # Parameters 496545 103690 103738 495298 495558 # Attention Layers 49 15 15 53 53 Attention Layer Dim. 32 24 24 32 32 Task Head Hidden Dim. 192 144 144 N/A (Linear) N/A (Linear) # Epochs 2000 200 400 500 50 Warmup Epoch Ratio 0.1 0.05 0.1 0.1 0.1 Batch Size 200 200 200 200 200 Initial Learning Rate 1e-7 1e-7 1e-7 1e-7 1e-7 Peak Learning Rate 5e-4 1e-3 1e-3 1e-3 1e-3 Final Learning Rate 1e-7 1e-7 1e-7 3e-4 1e-7 Betas (0.95, 0.98) (0.95, 0.98) (0.95, 0.98) (0.95, 0.98) (0.95, 0.98) Weight Decay Factor 0.5 0.3 0.3 3.0 0.3 Label Smoothing Factor N/A (Regression) 0.1 0.1 0.1 0.1 Residual Connection Scale $\alpha$ 0.2 0.4 0.4 0.2 0.2 Random Walk Encoding Type RRWP RRWP RRWP RRWP RRWP (D-)RRWP Random Walk Length 32 24 24 32 32 Random Regular Graph Degree 6 6 6 6 6 Random Edge Removal Rate 0.1 0.1 0.1 0.5 0.5
@@
Model Peptides-func Peptides-struct PascalVOC-SP COCO-SP roman-empire # Parameters 500074 498315 501493 499377 2075730 # Attention Layers 48 48 53 53 24 Attention Layer Dim. 32 32 32 32 96 Task Head Hidden Dim. 192 192 N/A (Linear) N/A (Linear) N/A (Linear) # Epochs 500 500 500 100 4000 Warmup Epoch Ratio 0.1 0.1 0.1 0.1 0.1 Batch Size 200 200 200 200 1* Initial Learning Rate 1e-7 1e-7 1e-7 1e-7 1e-7 Peak Learning Rate 1e-3 1e-3 1e-3 1e-3 1e-3 Final Learning Rate 1e-7 1e-7 1e-7 1e-7 1e-7 Betas (0.95, 0.98) (0.95, 0.98) (0.95, 0.98) (0.95, 0.98) (0.95, 0.98) Weight Decay Factor 0.3 3.0 1.0 0.3 1.0 Label Smoothing Factor 0.1 N/A (Regression) 0.1 0.1 0.1 Residual Connection Scale $\alpha$ 0.2 0.2 0.2 0.2 0.3 Random Walk Encoding Type D-RRWP RRWP D-RRWP D-RRWP D-RRWP (D-)RRWP Random Walk Length 128 64 128 64 256 Random Regular Graph Degree 3 3 6 6 3 Random Edge Removal Rate 0.1 0.1 0.1 0.1 0.5